OpenStudy (anonymous):

If you have an x-intercept = -2 and a y-intercept = -5 find the slope and the y-intercept

OpenStudy (jhannybean):

x-intercept means you have a point on the x-axis,this point is labeled (-2,0) the y-intercept is the same thing, but now you have a point on the y-axis, labeled (0,-5) In order to find the slope we will utilize the formula \[\large m=\frac{y_{2}-y_{1}}{x_{2}-x_{1}}\] where (x1,y1) = (-2,0) and (x2,y2) = (0,-5) \[\large m= \frac{-5-0}{0-(-2)}= -\frac{5}{2}\] and the y-intercept is given to us already.